A projectile is to be launched at an angle of 30 degrees so that it falls beyond a point of length 20m.

what is the range of value of the initial velocity so that the projectile falls between the points m and n


1
Expert's answer
2021-11-03T10:18:37-0400

The velocity is


"v=\sqrt{Rg/\sin(2 \theta)}=15\text{ m/s}."